# X-Fi: A Modality-Invariant Foundation Model for Multimodal Human Sensing

### v2

@@ Abstract
Human sensing, which employs various sensors and advanced deep learning technologies to accurately capture and interpret human body information, has significantly impacted fields like public security and robotics. However, current human sensing primarily depends on modalities such as cameras and LiDAR, each of which has its own strengths and limitations. Furthermore, existing multi-modal fusion solutions are typically designed for fixed modality combinations, requiring extensive retraining when modalities are added or removed for diverse scenarios. In this paper, we propose a modality-invariant foundation model for all modalities, X-Fi, to address this issue. X-Fi enables the independent or combinatory use of sensor modalities without additional training by utilizing a transformer structure to accommodate variable input sizes and incorporating a novel “X-fusion” mechanism to preserve modality-specific features during multimodal integration. This approach not only enhances adaptability but also facilitates the learning of complementary features across modalities. Extensive experiments conducted on the MM-Fi and XRF55 datasets, employing six distinct modalities, demonstrate that X-Fi achieves state-of-the-art performance in human pose estimation (HPE) and human activity recognition (HAR) tasks. The findings indicate that our proposed model can efficiently support a wide range of human sensing applications, ultimately contributing to the evolution of scalable, multimodal sensing technologies.
@@ 1 Introduction
Human sensing refers to using various sensors to capture human body information in a specific space, such as presence (Nguyen et al., 2016; Nanzer, 2017), activity (Vrigkas et al., 2015; Yang et al., 2023)), and pose (Toshev & Szegedy, 2014; Cai et al., 2022). Benefiting from advanced deep learning technology, human sensing has made significant progress in both accuracy and granularity, expanding applications in many domains. For instance, video surveillance (Elharrouss et al., 2021) is empowered by human sensing algorithms to analyze dangerous behaviors. In the field of robotics, human pose helps robots comprehend human actions and then realize human-robot collaboration (Gao et al., 2019).
Currently, human sensing tasks mainly rely on vision-based modalities like cameras (Ionescu et al., 2013), which face inherent limitations such as reliance on illumination, privacy concerns, and insufficient 3D information (Cai et al., 2022). Alternative modalities like LiDAR, mmWave radar, and WiFi have be introduced to overcome these challenges (Nirmal et al., 2021). Each modality has distinctive advantages but also has limitations, e.g., LiDAR offers high-resolution data but is expensive (Li et al., 2022), mmWava radar is cost-effective and informative but lacks sensitivity to static objects (An & Ogras, 2021), and WiFi is ubiquitous and privacy-preserving but has low resolution (Yang et al., 2023; Zhou et al., 2023). Therefore, single-modal sensing solutions cannot fit all scenarios, and a multi-modal approach that leverages the strengths of each modality is essential for future advancements in human sensing.
@@
Figure 1: The left image shows existing modality-corresponding human sensing solutions, while the right image illustrates our proposed modality-invariant foundation model, X-Fi, which enables modality-invariant human sensing.
@@
To achieve modality-invariant human sensing, we aim to design a foundation model that can dynamically handle varying numbers of modalities. First, considering the arbitrary number of input modalities, we naturally resorted to a transformer structure due to its adaptability to variable input sizes. Second, recognizing that each modality has unique characteristics, we adopt modality-specific feature extractors to capture distinctive information. Finally, to retain each modality’s distinct features during multimodal fusion, we introduced a cross-attention fusion mechanism that reduces the information loss caused by multi-modal feature compression and fusion process. Surprisingly, this design not only supports dynamic inputs from various modalities but by training with diverse modality combinations to perform sensing tasks, it learns complementary features among modalities as well. As a result, it even achieves superior performance against conventional modality fusion methods such as feature-level and decision-level fusion.
@@
To this end, we propose a novel modality-invariant foundation model, X-Fi, for versatile human sensing. X-Fi can take in any combinations of modalities and activate corresponding parts to extract modality-specific features. A cross-modal transformer is designed to learn a cross-modal feature and each modality information is then preserved in the representation by executing cross-attention multi-modal fusion processes to preserve distinctive modal features. As shown in Figure.2, we have a set of modality-corresponded feature encoders and key-value generators to extract, fine-tune and preserve modality-specific features. For versatile combinations of sensor modalities, corresponding feature encoders are activated to obtain modality-specific features of each modality. Subsequently, the attention-based key-value generators are trained to produce key-value pairs, which contain more task-related information derived from modality-specific features (Vaswani et al., 2017). For the modality-invariant multi-modal fusion process, we introduce “X-fusion” block. It consists of a cross-modal transformer and a set of modality-correlated cross-attention transformers. The cross-modal transformer is designed to adaptively learn a unified embedding across variable-sized multi-modal embedding, which comprises the features of each modality. The generated cross-modal embedding serves as the query, along with key-value pairs of each modality, are sent to the corresponding cross-attention transformer to perform attention-based feature injection. We evaluated X-Fi on human pose estimation (HPE) and human activity recognition (HAR) tasks in MM-Fi (Yang et al., 2024) and XRF55 (Wang et al., 2024), demonstrated that X-Fi surpasses previous methods by MPJPE 24.8% and PA-MPJPE 21.4% on HPE task, 2.8% on HAR task.
@@ 2.1 Human Sensing
Human sensing has been extensively explored for applications such as autonomous driving and robotics in recent years, aiming to capture, analyze, and interpret human body information through various sensors and novel deep learning frameworks (Cai et al., 2022). Most human sensing approaches could generally be split into three catgories: i) Vision-based approaches, where human body information is extracted from visual data captured by cameras (Shahroudy et al., 2016) and leveraged by deep learning architectures like convolutional structures (Chun et al., 2018; Pavlakos et al., 2018; Sun et al., 2019; Li et al., 2020) and transformer (Goel et al., 2023; Cai et al., 2024; Qiu et al., 2022). ii) Sensor-based approaches, using alternative modalities such as LiDAR (Ren et al., 2024), mmWave radar (Wang et al., 2021; An & Ogras, 2022), and WiFi CSI (Yang et al., 2023; Zhou et al., 2023) to overcome limitations of vision-based methods and expand applicability. iii) Multimodal fusion approaches, including decision-level fusion (An et al., 2022; Yang et al., 2024) and feature-level fusion (Zheng et al., 2022; Chen et al., 2023), which combine each modality’s strengths to improve sensing performance.
@@ 2.2 Modality-Invariant Methods
Current human sensing methods lack broad applicability. Both decision-level fusion approaches (Yang et al., 2024) and feature-level fusion approaches (Zheng et al., 2022; Chen et al., 2023) are typically tailored to fixed modality combinations for specific applications. For example, (Zheng et al., 2022) developed a specific heatmap sampling structure to fuse 2D heatmaps from RGB images with projected 2D point clouds from LiDAR. Similarly, (Chen et al., 2023) employed modality-specific backbones to extract grid features from images and cluster features from mmWave point clouds, combining them via a global integration module to produce fused features.
In fields beyond human sensing, several studies on multimodal fusion seek to accommodate diverse sensor inputs to a unified model for various applications. Some methods align all input modalities with a primary modality (Man et al., 2023; Han et al., 2024), like OneLLM, which employs a universal encoder and projection module to tokenize eight modalities and align the resulting multimodal tokens with language using large language models (Han et al., 2024), but these approaches heavily rely on the presence of the primary modality. Some other methods, such as (Xue & Marculescu, 2023), design a gating network to identify the active input modalities and subsequently activate the modality-combination-specific expert networks for multimodal fusion. Yet, this approach is struggled by the necessity of creating numerous expert networks to accommodate a wide range of potential modal combinations in the presence of an extensive sensor options. Other methods like (Memmel et al., 2023) utilize a transformer-based structure to process tokenized input and also introduce a modality-invariant training strategy by dropping modalities during training. However, this approach is only applicable when the input modalities are limited to vision-based modalities.
Figure 2: The architecture of the proposed modality-invariant foundation model, X-Fi. The X-Fi consists of a modality feature encoding module that extracts modality-specific features, and a X-Fusion modules that learns an unified cross-modal feature through modality-fusion in the Cross Modal Transformer and preserves distinctive modal features through feature injection in modality-specified Cross Attention Transformers.
@@ 3 Modality-Invariant Foundation Model for Human Sensing
Assume that input $X$ comprises a random combination of modalities within a specific range, including any number and type of modalities. Our objective is to learn a unified representation $Emb_{cm}$ that integrates modality-specific features from $X$ by using a modality-invariant foundation model $\mathcal{F}$. This representation can be mapped to various down-stream human sensing task spaces by using corresponding task-specific functions $g_{t_{i}}$, such that the task-specified loss function $\mathcal{L}_{t_{i}}(g_{t_{i}}(\mathcal{F}(X));y_{t_{i}})$ is minimized. Thus we propose X-Fi, a modality-invariant human sensing framework designed to map the arbitrary multimodal input $X$ to any task-specific target domain $t_{i}$, such as HPE and HAR.
As illustrated in Fig 2, given $N$ modalities in the dataset, we first prepare a set of modality-specified feature encoders ${\left\{Enc_{i}\left(\cdot;\theta_{Enc,i}\right)\right\}}_{i=1}^{N}$ to extract features ${\left\{F_{i}\in\mathbb{R}^{n_{f}\times d_{f}}\right\}}_{i=1}^{N}$ form corresponding modality inputs ${\left\{x_{i}\right\}}_{i=1}^{N}$. Then we introduce cross-attention-based multi-modal fusion design that could efficiently preserve modality-specific features in the learned unified cross-model embedding. The modality-specified key-value generators ${\left\{G_{kv,i}\left(\cdot;\theta_{G,i}^{kv}\right)\right\}}_{i=1}^{N}$ are applied to encode the modality information from ${\left\{F_{i}\right\}}_{i=1}^{N}$ into key-value pairs ${\left\{(K_{i},V_{i})\right\}}_{i=1}^{N}$. To obtain an unified cross modal embedding $Emb_{cm}\in\mathbb{R}^{n_{f}\times d_{f}}$, we use a cross modal transformer $G_{cm}\left(\cdot;\theta_{G}^{cm}\right)$ to learn from the $Emb_{mm}\in\mathbb{R}^{(n_{f}*N)\times d_{f}}$. The generated cross modal embedding $Emb_{cm}$ serves as the query $Q$, together with each key-value pair, to send into the modality-corresponded cross attention transformer ${\left\{G_{ca,i}\left(\cdot;\theta_{G,i}^{ca}\right)\right\}}_{i=1}^{N}$ in order to inject modality-specific information. Through iterative learning of the $Emb_{cm}$ and performing modality-specific information injection, the final produced $Emb_{cm}$ could be used for various downstream tasks via task-correlated Multi-layer Perceptron (MLP) head.
@@ 3.1 Modality Feature Encoding
In human sensing, different modalities exhibit distinctive characteristics, requiring unique feature encoder structures to extract robust modality-specific features. For instance, images are represented as a 2D matrix and have rich spatial feature, requiring models like CNN (LeCun et al., 1998) and ViT (Dosovitskiy et al., 2021) to extract vision features. While point clouds from LiDAR consist of a set of unordered points in Euclidean space, necessitating point cloud-specific frameworks like Point Net (Qi et al., 2017) and Point Transformer (Zhao et al., 2021) to process them. Thus, we introduce modality-specific feature encoders in X-Fi to extract features from various modalities.
In X-Fi, the modality-specified feature encoder $Enc_{i}\left(\cdot;\theta_{Enc,i}\right)$ for each modality consists of a pretrained feature extractor $E_{i}\left(\cdot;\theta_{E,i}\right)$ and a linear projection head. The feature extractors are responsible for extracting informative high-dimensional representations from modal data. To accommodate diverse combinations of sensor modalities across different application scenarios, we can select any model structures as backbones for these feature extractors. Each feature extractor is modality specified, which allows us to choose a model structure that best aligns with the unique characteristics of the modality’s data, thereby ensuring the efficiency of feature extraction. For example: for point clouds from mmWave radar and LiDAR, we adopt the Point Transformer backbone (Zhao et al., 2021). The self-attention mechanism in Point Transformers is particularly suitable for point cloud data due to its permutation invariance and ability to dynamically focus on relevant points.
Since the data structures of the features extracted by different modality-specific feature extractors are still different across modalities, we use a set of linear projection structures to unify these features into the same dimensional space,allowing them to be used together in subsequent stages of the model. Finally, the positional encoding is added to obtain the final features.
$\left\{Enc_{i}(\cdot;\theta_{Enc,i})\right\}_{i=1}^{N}=\left\{\text{LinearProjection}_{i}(E_{i}(\cdot;\theta_{E,i});\theta_{LP,i})\right\}_{i=1}^{N}$ (1)
@@ 3.2 Attention-based X-Fusion
After obtaining the modality-specific features, a natural question is posed: How can the modality-invariant foundation model effectively combine these features to obtain a unified cross-modal representation that is applicable for downstream human sensing tasks? Thus, in this section, we will interpret the proposed X-Fusion structure which consists of a cross-modal transformer to learn an unified cross-modal feature and cross-attention multi-modal fusion modules to preserve distinctive modality features.
@@
First, we concatenate all obtained modality-specific features to initialize a multi-modal embedding $Emb_{mm}\in\mathbb{R}^{(n_{f}*N)\times d_{f}}$. As illustrated in the Figure. 2, we then linearly map the multi-modal embedding into three vectors, named $Q_{mm},K_{mm},V_{mm}$. Next, we perform dot-product-based multi-head attention to calculate the attention value which models the dependencies among modality features.
Notably, The multi-head attention and feed-forward blocks output embedding of the same size as the input, resulting in variable-sized cross-modal embedding based on the number of modality inputs. This variability complicates the design of subsequent fusion blocks and task-specific projection heads which have fixed-size parameters. To address this issue, we apply adaptive pooling both after the multi-head attention block and within the residual connection to perform down sampling and obtain unified-sized outputs $Z_{mm}\in\mathbb{R}^{n_{f}\times d_{f}}$. Finally, we feed $Z_{mm}$ into a feed-forward structure with a residual connection to obtain the unified cross-modal embedding $Emb_{cm}\in\mathbb{R}^{n_{f}\times d_{f}}$.
$Z_{mm}=\text{AvgPool}(\text{MultiHead}\left(Q_{mm},K_{mm},V_{mm}\right))+\text{AvgPool}(Emb_{mm})$ (2)
$Emb_{cm}=FFN(Z_{mm})+Z_{mm}$ (3)
@@
However, the basic transformer design, which has a layered architecture with sequentially stacked transformer blocks, often leads to a prioritization of dominant modalities. First reason is that the self-attention mechanism in the cross-modal transformer tends to assign higher attention weights to the more informative modality-specific features (Lu et al., 2019), causing significant information loss from less dominant modalities. Secondly, in layered structures, different layers focus on distinct aspects of the input data, and later layers naturally exerting a greater influence on the overall output. As a result, the modalities emphasized by the later layers tend to be disproportionately prioritized. To efficiently and equally fuse features from any combination of modalities, we develop a cross-attention multi-modal fusion architecture that ensures the model pays equal attention to every input modality by incorporating modality-specific information into the cross-modal embedding. We also design an X-Fusion process that iterates on the same block, compelling the X-Fusion block to identify a unified dynamic attention allocation method that effectively preserves the features of each modality and optimally integrates them to achieve the most optimal solution.
@@
To prepare key-value pairs that preserve modality-specific information, we design key-value generators for each modality. take the i-th modality for illustration, key-value generator $G_{kv,i}$ receives modality specific feature $F_{i}\in\mathbb{R}^{n_{f}\times d_{f}}$ extracted from feature extractor $E_{i}$ as the input. $F_{i}$ is first passed through a two-layer MLP with a hidden dimension of $d_{f}*2$ incorporating a Rectified Linear Unit (ReLU) activation (Nair & Hinton, 2010) and layer normalization in between. The MLP output is then projected to key $K_{i}$ and value $V_{i}$ using separate linear mapping layers. Once the key-value pair $(K_{i},V_{i})$ is generated from $F_{i}$, it remains unchanged during iterations in the X-fusion block to prevent information leakage between modalities.
Cross-Attention Transformers.
We employ a series of cross-attention transformers to integrate modality-specific information into the cross-modal embedding $Emb_{cm}$. Each transformer is associated with a specific modality and receives three inputs: the unified cross-modal embedding $Emb_{cm}$ that is shared across all cross-attention transformers, the key ($K$) and value ($V$) pairs obtained from modality-specific key-value generators. Take the i-th modality as an example, $Emb_{cm}$ serves as the query ($Q$) in the multi-head attention mechanism. Attention weights are computed based on $Q$ and $K_{i}$ to scale the modality-specific value $V_{i}$, yielding an intermediate result denoted as $O_{i}$. Subsequently, $Emb_{cm}$ is residual added to $O_{i}$ to retain cross-modal features, resulting in $O_{i}^{{}^{\prime}}$. The final output $F^{\prime}_{i}$ of each modality-specific cross-attention transformer is produced by passing $O_{i}^{{}^{\prime}}$ through a feed-forward structure with a residual connection. Each $F^{\prime}_{i}$ generated from modality-specific cross-attention transformer encapsulates fused information focusing on a single modality. To integrate information from all modalities, we concatenate all ${\left\{F^{\prime}_{i}\right\}}_{i=1}^{N}$ to get the updated multi-modal embedding $Emb_{mm}^{{}^{\prime}}$, which is used in the subsequent X-fusion block iteration.
$Emb_{mm}^{{}^{\prime}}=\text{Concat}{\left\{G_{ca,i}\left(Emb_{cm},K_{i},V_{i};\theta_{G,i}^{ca}\right)\right\}}_{i=1}^{N}$ (4)

### v3

@@ Abstract
Human sensing, which employs various sensors and advanced deep learning technologies to accurately capture and interpret human body information, has significantly impacted fields like public security and robotics. However, current human sensing primarily depends on modalities such as cameras and LiDAR, each of which has its own strengths and limitations. Furthermore, existing multimodal fusion solutions are typically designed for fixed modality combinations, requiring extensive retraining when modalities are added or removed for diverse scenarios. In this paper, we propose a modality-invariant foundation model for all modalities, X-Fi, to address these issues. X-Fi enables the independent or combinatory use of sensor modalities without additional training by utilizing a transformer structure to accommodate variable input sizes and incorporating a novel “X-fusion” mechanism to preserve modality-specific features during multimodal integration. This approach not only enhances adaptability but also facilitates the learning of complementary features across modalities. Extensive experiments conducted on the MM-Fi and XRF55 datasets, employing six distinct modalities, demonstrate that X-Fi achieves state-of-the-art performance in human pose estimation (HPE) and human activity recognition (HAR) tasks. The findings indicate that our proposed model can efficiently support a wide range of human sensing applications, ultimately contributing to the evolution of scalable, multimodal sensing technologies.
@@ 1 Introduction
Human sensing refers to using various sensors to capture human body information in a specific space, such as presence (Nguyen et al., 2016; Nanzer, 2017), activity (Vrigkas et al., 2015; Yang et al., 2023)), and pose (Toshev & Szegedy, 2014; Cai et al., 2022). Benefiting from advanced deep learning technology, human sensing has made significant progress in both accuracy and granularity, expanding applications in many domains. For instance, video surveillance (Elharrouss et al., 2021) is empowered by human sensing algorithms to analyze dangerous behaviors. In the field of robotics, human sensing helps robots comprehend human actions and realize human-robot collaboration (Gao et al., 2019).
Currently, human sensing tasks mainly rely on vision-based modalities like cameras (Ionescu et al., 2013), which face inherent limitations such as reliance on illumination, privacy concerns, and insufficient 3D information (Cai et al., 2022). Alternative modalities like LiDAR, mmWave radar, and WiFi have been introduced to overcome these challenges (Nirmal et al., 2021). Each modality has distinctive advantages but also has limitations, e.g., LiDAR offers high-resolution data but is expensive (Li et al., 2022), mmWava radar is cost-effective and informative but lacks sensitivity to static objects (An & Ogras, 2021), and WiFi is ubiquitous and privacy-preserving but has low resolution (Yang et al., 2023; Zhou et al., 2023). Therefore, single-modal sensing solutions cannot fit all scenarios, and a multi-modal approach that leverages the strengths of each modality is essential for future advancements in human sensing.
@@
Figure 1: The left image depicts current human sensing solutions that are specifically designed and trained for fixed modality combinations, while the right image illustrates our proposed modality-invariant foundation model, X-Fi, which can be trained once and adapted to various scenarios.
@@
To achieve modality-invariant human sensing, we aim to design a foundation model that can dynamically handle varying numbers of modalities. First, considering the arbitrary number of input modalities, we naturally resort to a transformer structure due to its adaptability to variable input sizes (Vaswani et al., 2017). Second, recognizing that each modality has unique characteristics, we adopt modality-specific feature extractors to capture distinctive information. Finally, to retain each modality’s distinct features during multimodal fusion, we introduce a cross-attention fusion mechanism that reduces the information loss caused by multimodal feature compression and fusion process. Surprisingly, this design not only supports dynamic inputs from various modalities but by training with diverse modality combinations to perform sensing tasks, it learns complementary features among modalities as well. As a result, it even achieves superior performance against conventional modality fusion methods such as feature-level and decision-level fusion.
@@
To this end, we propose a novel modality-invariant foundation model, X-Fi, for versatile human sensing. X-Fi can take in any combination of modalities and activate corresponding parts to extract modality-specific features. A cross-modal transformer is designed to learn a cross-modal feature and each modality information is then preserved in the representation by executing cross-attention multi-modal fusion processes to preserve distinctive modal features. As shown in the Figure 2, we have a set of modality-corresponded feature encoders and key-value generators to extract, fine-tune and preserve modality-specific features. For versatile combinations of sensor modalities, corresponding feature encoders are activated to obtain modality-specific features of each modality. Subsequently, the key-value generators are trained to produce key-value pairs, which contain more task-related information derived from modality-specific features. For the modality-invariant multimodal fusion process, we introduce “X-fusion” block. It consists of a cross-modal transformer and a set of modality-correlated cross-attention modules. The cross-modal transformer is designed to adaptively learn a unified cross-modal embedding across variable-sized multi-modal embedding, which comprises the features of each modality. The generated cross-modal embedding serves as the query, along with key-value pairs of each modality, are sent to the corresponding cross-attention module to perform attention-based feature injection. We evaluated X-Fi on human pose estimation (HPE) and human activity recognition (HAR) tasks on MM-Fi (Yang et al., 2024) and XRF55 (Wang et al., 2024) datasets, demonstrated that X-Fi surpasses previous methods by MPJPE 24.8% and PA-MPJPE 21.4% on HPE task, 2.8% on HAR task.
@@ 2.1 Human Sensing
Human sensing has been extensively explored for applications such as autonomous driving and robotics in recent years, aiming to capture, analyze, and interpret human body information through various sensors and novel deep learning frameworks (Cai et al., 2022). Human sensing approaches are typically divided into three categories: i) Vision-based approaches, where human body information is extracted from visual data captured by cameras (Shahroudy et al., 2016) and leveraged by deep learning architectures like convolutional structures (Pavlakos et al., 2018; Sun et al., 2019; Li et al., 2020) and transformer (Goel et al., 2023; Cai et al., 2024). ii) Sensor-based approaches, using alternative modalities such as LiDAR (Ren et al., 2024), mmWave radar (Wang et al., 2021; An & Ogras, 2022), and WiFi CSI (Yang et al., 2023; Zhou et al., 2023) to overcome limitations of vision-based methods and expand applicability. iii) Multimodal fusion approaches, including decision-level fusion (An et al., 2022; Yang et al., 2024) and feature-level fusion (Zheng et al., 2022; Chen et al., 2023), which combine each modality’s strengths to improve sensing performance.
@@ 2.2 Modality-Invariant Methods
Current human sensing methods lack broad applicability. Feature-level fusion approaches (Zheng et al., 2022; Chen et al., 2023) are typically tailored to fixed modality combinations for specific applications. For example, (Chen et al., 2023) employed modality-specific backbones to extract grid features from images and cluster features from mmWave point clouds, combining them via a global integration module to produce fused features. Decision-level fusion approaches (Yang et al., 2024; Yadav et al., 2023) are easy to adapt to various modality combinations, but the performance could be hindered by less informative modalities, e.g. WiFi.
In fields beyond human sensing, several studies on multimodal fusion seek to accommodate diverse sensor inputs to a unified model for various applications. Some approaches apply cross-modal distillation techniques to seek the shared and the supplementary information across multi-modalities (Xue et al., 2022; Kong et al., 2019; Wang et al., 2020; Bruce et al., 2021). Some methods align multi-modalities features in a joint embedding space with a unified model structure (Man et al., 2023; Radford et al., 2021; Driess et al., 2023; Zhang et al., 2023; Girdhar et al., 2023; Moon et al., 2023; Tong et al., 2021; Wu et al., 2023). However, these approaches require retraining or fine-tuning to adapt to various scenarios. In recent works, foundation models, which are large-scale, pre-trained models that learn unified representations from diverse inputs, have gained attention for their capacity to be fine-tuned for a wide range of tasks (Devlin, 2018; Touvron et al., 2023). Extending this concept, modality-invariant representation learning have been proposed to handle dynamic multimodal combinations across various scenarios.  (Xue & Marculescu, 2023) designs a gating network to dynamically activate the modality-combination-specific expert networks for multimodal fusion and  (Memmel et al., 2023) introduces a modality-invariant training strategy by dropping modalities during training. However, these approaches face challenges due to the need for numerous networks and configurations to handle diverse modality combinations in the presence of extensive sensor options.
Figure 2: The architecture of the proposed modality-invariant foundation model, X-Fi. X-Fi consists modality feature encoders and an X-Fusion module, which includes a cross-modal transformer and modality-specified cross-attention modules. The modalities with dotted lines represent inactivate modalities in the given scenario. The $N$ in X-Fusion block represents the number of iterations.
@@ 3 Modality-Invariant Foundation Model for Human Sensing
Assume the input $X$ comprises a random combination of modalities within a specific range, including any number and type of modalities. Our objective is to learn a unified representation $z_{cm}$ that integrates modality-specific features from $X$ by using a modality-invariant foundation model $\mathcal{F}$. This representation can be mapped to various down-stream human sensing task spaces $t_{i}$ by using corresponding task-specific functions $g_{t_{i}}$, such that the task-specified loss function $\mathcal{L}_{t_{i}}(g_{t_{i}}(\mathcal{F}(X));y_{t_{i}})$ is minimized. In this paper, the $y_{t_{i}}$ refers to 3D human joints for HPE task and human activity classes for HAR task. Thus we propose X-Fi, a modality-invariant human sensing framework designed to map the arbitrary multimodal input $X$ to any task-specific target domain $t_{i}$.
As illustrated in the Figure 2, given $k$ modalities in the dataset, we first prepare a set of modality-specified feature encoders ${\left\{E_{i}\left(\cdot;\theta_{E,i}\right)\right\}}_{i=1}^{k}$ to extract features ${\left\{f_{i}\in\mathbb{R}^{n_{f}\times d_{f}}\right\}}_{i=1}^{k}$ from corresponding modality inputs ${\left\{x_{i}\right\}}_{i=1}^{k}$, where $n_{f}$ indicates the number of features extracted for each modality and $d_{f}$ indicates the dimension of each feature. Then we introduce a cross-attention-based multi-modal fusion design that could efficiently preserve modality-specific features in the learned unified cross-model embedding. The modality-specified key-value generators ${\left\{G_{kv,i}\left(\cdot;\theta_{G,i}^{kv}\right)\right\}}_{i=1}^{k}$ are applied to encode the modality information from ${\left\{f_{i}\right\}}_{i=1}^{k}$ into key-value pairs ${\left\{(K_{i},V_{i})\right\}}_{i=1}^{k}$. To obtain an unified cross-modal embedding $z_{cm}\in\mathbb{R}^{n_{f}\times d_{f}}$, we use a cross-modal transformer $G_{cm}\left(\cdot;\theta_{G}^{cm}\right)$ to learn from the $z_{mm}\in\mathbb{R}^{(n_{f}*k)\times d_{f}}$. The generated cross modal embedding $z_{cm}$ serves as the query $Q$, together with each key-value pair, to send into the modality-specific cross-attention modules ${\left\{G_{ca,i}\left(\cdot;\theta_{G,i}^{ca}\right)\right\}}_{i=1}^{k}$ in order to inject modality-specific information. Through iterative learning of the $z_{cm}$ and performing modality-specific information injection, the final produced $z_{cm}$ could be used for various downstream tasks via task-correlated Multi-layer Perception (MLP) head.
@@ 3.1 Modality Feature Encoding
In human sensing, different modalities exhibit distinctive characteristics, requiring unique feature encoder structures to extract robust modality-specific feature. For instance, images are represented as a 2D matrices and have rich spatial features, requiring models like CNN (LeCun et al., 1998) and ViT (Dosovitskiy, 2020) to extract vision features. While point clouds from LiDAR consist of a set of unordered points in Euclidean space, necessitating point cloud-specific frameworks like Point Net (Qi et al., 2017) and Point Transformer (Zhao et al., 2021) to process them. Thus, we introduce modality-specific feature encoders in X-Fi to extract features from various modalities.
In X-Fi, the modality-specified feature encoder $E_{i}\left(\cdot;\theta_{E,i}\right)$ for each modality consists of a pretrained feature extractor $F_{i}\left(\cdot;\theta_{F,i}\right)$ and a linear projection head. The feature extractors are responsible for extracting informative high-dimensional representations from modal data. To accommodate diverse combinations of sensor modalities across different application scenarios, we can select any model structures as backbones for these feature extractors. Each feature extractor is modality specified, which allows us to choose a model structure that best aligns with the unique characteristics of the modality’s data, thereby ensuring the efficiency of feature extraction. For example: for point clouds from mmWave radar and LiDAR, we adopt the Point Transformer backbone (Zhao et al., 2021). The self-attention mechanism in Point Transformers is particularly suitable for point cloud data due to its permutation invariance and ability to dynamically focus on relevant points.
Since the data structures of the features extracted by different modality-specific feature extractors are still different across modalities, we use a set of linear projection structures to unify these features into the same dimensional space, allowing them to be used together in subsequent stages of the model. Finally, positional encoding is added to obtain the final features.
$\left\{E_{i}(\cdot;\theta_{E,i})\right\}_{i=1}^{k}=\left\{\text{LinearProjection}_{i}(F_{i}(\cdot;\theta_{F,i});\theta_{LP,i})\right\}_{i=1}^{k}$ (1)
@@ 3.2 Attention-based X-Fusion
After obtaining the modality-specific features, a natural question is posed: How can the modality-invariant foundation model effectively combine these features to obtain a unified cross-modal representation that is applicable for downstream human sensing tasks? Thus, in this section, we will interpret the proposed X-Fusion structure which consists of a cross-modal transformer to learn unified cross-modal features and cross-attention fusion modules to preserve distinctive modality features.
@@
First, we concatenate all obtained modality-specific features to initialize a multi-modal embedding $z_{mm}\in\mathbb{R}^{(n_{f}*k)\times d_{f}}$. As illustrated in the Figure 2, we then linearly map the multi-modal embedding into three vectors, named $Q_{mm},K_{mm},V_{mm}$. Next, we perform dot-product-based multi-head attention to calculate the attention value which models the dependencies among modality features.
Notably, the multi-head attention and feed-forward blocks output embedding of the same size as the input, resulting in variable-sized cross-modal embedding based on the number of modality inputs. This variability complicates the design of subsequent fusion blocks and task-specific projection heads which have fixed-size parameters. To address this issue, we apply adaptive pooling both after the multi-head attention block and within the residual connection to perform downsampling and obtain unified-sized outputs $o_{mm}\in\mathbb{R}^{n_{f}\times d_{f}}$. Finally, we feed $o_{mm}$ into a feed-forward structure with a residual connection to obtain the unified cross-modal embedding $z_{cm}\in\mathbb{R}^{n_{f}\times d_{f}}$.
$o_{mm}=\text{AvgPool}(\text{MultiHead}\left(Q_{mm},K_{mm},V_{mm}\right))+\text{AvgPool}(z_{mm})$ (2)
$z_{cm}=FFN(o_{mm})+o_{mm}$ (3)
The adaptive average pooling performs the dimension reduction for the multimodal feature and the sequential feature, which are summed up for the following layer. Then the random modality dropping in training ensures that all possible combinations of modalities are learned by X-Fi, implicitly aligning the feature of each modality. Therefore the final feature can be modality-invariant, demonstrated by massive empirical studies using 38 modality combinations on two large-scale datasets.
@@
However, the basic transformer design, which has a layered architecture with sequentially stacked transformer blocks, often prioritize certain modalities during training and the inference performance will rely on the existence of the dominant modalities. The reason is that the self-attention mechanism in the cross-modal module tends to focus on the more intuitive and informative modality-specific features during training (Lu et al., 2019). Such focus results in significant information loss from less dominant modalities, compromising robust and equitable performance across different scenarios where dominant modalities may be distorted or missing (Ma et al., 2022; Memmel et al., 2023). To address this, we propose cross-attention fusion modules that reintegrate overlooked modality-specific information into the cross-modal embeddings, thereby encouraging effective learning from less-informative modalities during training. Rather than impeding the features extraction from more informative modalities, such approach enhances the model’s ability to capture complementary features across modalities, improving robustness across diverse application scenarios. We also design an X-Fusion process that iterates on the same block, compelling the X-Fusion block to identify a unified dynamic attention allocation method that effectively preserves the features of each modality and optimally integrates them to achieve the most optimal solution.
@@
To prepare key-value pairs that preserve modality-specific information, we design key-value generators for each modality. As illustrated in the Figure 7, key-value generator $G_{kv,i}$ receives modality specific feature $f_{i}\in\mathbb{R}^{n_{f}\times d_{f}}$ extracted from feature encoder $E_{i}$ as the input. $f_{i}$ is first passed through a two-layer MLP with a hidden dimension of $d_{f}*2$ incorporating a Rectified Linear Unit (ReLU) activation (Nair & Hinton, 2010) and layer normalization in between. The MLP output is then projected to key $K_{i}$ and value $V_{i}$ using separate linear mapping layers. Once the key-value pair $(K_{i},V_{i})$ is generated from $f_{i}$, it remains unchanged during iterations in the X-fusion block to prevent information leakage between modalities.
Cross-Attention Modules.
We employ a series of cross-attention modules to integrate modality-specific information into the cross-modal embedding $z_{cm}$. Each transformer is associated with a specific modality and receives three inputs: the unified cross-modal embedding $z_{cm}$ that is shared across all cross-attention modules, the key ($K$) and value ($V$) pairs obtained from modality-specific key-value generators. Take the i-th modality as an example, as illustrated in the Figure 7, $z_{cm}$ serves as the query ($Q$) in the multi-head attention mechanism. Attention weights are computed based on $Q$ and $K_{i}$ to scale the modality-specific value $V_{i}$, yielding an intermediate result denoted as $o_{ca,i}$. Subsequently, $z_{cm}$ is residual added to $o_{ca,i}$ to retain cross-modal features, resulting in $o_{ca,i}^{{}^{\prime}}$. The final output $f^{\prime}_{i}$ of each modality-specific cross-attention module is produced by passing $o_{ca,i}^{{}^{\prime}}$ through a feed-forward structure with a residual connection. Each $f^{\prime}_{i}$ generated from the modality-specific cross-attention module encapsulates fused information focusing on a single modality. To integrate information from all modalities, we concatenate all ${\left\{f^{\prime}_{i}\right\}}_{i=1}^{k}$ to get the updated multi-modal embedding $z_{mm}^{{}^{\prime}}$, which is used in the subsequent X-fusion block iteration.
$z_{mm}^{{}^{\prime}}=\text{Concat}{\left\{G_{ca,i}\left(z_{cm},K_{i},V_{i};\theta_{G,i}^{ca}\right)\right\}}_{i=1}^{k}$ (4)